OpenStudy (chrispaul03):

The sum of two numbers is 60. The greater number is 6 more than the smaller number. Which equation can be used to solve for the smaller number? x + (x + 6) = 60 x − (x + 6) = 60 x(x + 6) = 60 x(x− 6) = 60 I need help please

OpenStudy (anonymous):

A

OpenStudy (anonymous):

its not

OpenStudy (anonymous):

The first one x + (x + 6) = 60

OpenStudy (anonymous):

no

OpenStudy (anonymous):

because you would need 34(34+6)=74

OpenStudy (anonymous):

The smaller number is x and the bigger number is 6 more than the smaller number, hence x + 6. Together they equal 60

OpenStudy (anonymous):

I honestly don't know, you might be right, but that's what I think the answer is

OpenStudy (anonymous):

Im stumped too

OpenStudy (chrispaul03):

so its a?

OpenStudy (anonymous):

I was the top student in Algebra and my end grade was 103%

OpenStudy (anonymous):

Yes I'm pretty sure it is A

OpenStudy (anonymous):

im not sure but

OpenStudy (anonymous):

If you solved it, x would be 27. 27 + (27 + 6) = 60 27 + 33 = 60

OpenStudy (anonymous):

Yes that works perfectly

OpenStudy (chrispaul03):

I have two more can you help me with them

OpenStudy (anonymous):

sure

OpenStudy (chrispaul03):

negative 7 over 6, multiplied by x minus 6 equals negative 48. 36 49 −36 −49

OpenStudy (anonymous):